The integrating factor for the Bernouli differential equation 2xyy' = y^2 - 2x^3 given initial condition y(1)=2


1
Expert's answer
2021-11-03T18:53:05-0400

Rewrite in the form of a first order ODE


"y'-\dfrac{1}{2x}y=-x^2y^{-1}"

Substitute "z=y^{1-(-1)}=y^2"


"z'=2yy'"

"2yy'-\dfrac{1}{x}y^2=-2x^2"

"z'-\dfrac{1}{x}z=-2x^2"

Integrating factor

"IF=\mu(x)=e^{\int(-1/x)dx}=\dfrac{1}{x}"

"\dfrac{1}{x}z'-\dfrac{1}{x}(\dfrac{1}{x})z=-2x^2(\dfrac{1}{x})"

"d(\dfrac{z}{x})=-2xdx"

Integrate


"\int d(\dfrac{z}{x})=-\int 2xdx"

"\dfrac{z}{x}=-x^2+C"

"z=-x^3+Cx"

"y^2=-x^3+Cx"

"y=\pm\sqrt{-x^3+Cx}"

Given initial condition "y(1)=2"


"2=\sqrt{-(1)^3+C(1)}"

"C=9"

"y=\sqrt{-x^3+9x}"